# 2Business Informatics Group, TU Wien, Austria

### 19 Dec 2023 v2

@@ 1 Introduction
The steadily accelerating innovation pathways of humankind have rendered prevailing systems engineering paradigms unsustainable. By Brundtland’s classic definition of sustainability (Brundtland, 1987), systems engineering falls short of “meeting the needs of the present without compromising the ability of future generations to meet their own needs”. Our systems engineering practices fail to fulfill the four essential sustainability dimensions of technical systems Penzenstadler and Femmer (2013): technical (long-term usage); economic (financial viability); environmental (reduced impact); and social (elevated utility).
@@
Clearly, systems engineering must change course. While most analysts agree that changing course will require major political and industry intervention Freitag et al. (2021), and some argue that fulfilling these goals is not feasible due to the limitations of our current frame of thinking Becker (2023), it is also expected that, in the decade ahead of us, users and organizations will reward and demand efforts toward sustainability. The importance of reuse and repurposing will increase and become a key driving force in systems engineering. Sustainability as a system characteristic will become a leading principle in systems’ design, operation, maintenance, and post-life. These trends have been collectively identified in the Systems Engineering Vision 2035 report International Council on Systems Engineering (2021) (INCOSE) of the International Council on Systems Engineering (INCOSE), the leading systems engineering society, as the top “global megatrend” that will instigate the development of radically new systems engineering frameworks, methods, and tools. The European Commission has also identified sustainability as a critical enabler of a more resilient European industry within the framework of Industry 5.0 European Commission (2020). Expert voices call for immediate action in devising frameworks, methods, and tools for sustainable systems engineering practices van der Aalst et al. (2023) and fostering a circular economy Kristoffersen et al. (2020).
@@
The rest of this article is structured as follows. In Section 2, we provide an overview of the background related to our work. In Section 3, we offer a definition of circular systems engineering. In Section 4, we discuss the foundational principles of circular systems engineering. In Section 5, we discuss sustainability maturity levels and strategies to reach the level of circularity in systems engineering. In Section 6, we outline the key challenges and research opportunities for the modeling community in implementing circular systems engineering. In Section 7, we wrap up the paper by drawing the conclusions.
@@ 2.2 Circular economy
Circular economy (CE) is a model of production and consumption in an economic system that promotes retaining value over numerous cycles of the system as long as possible. CE is one among several solutions for fostering sustainable systems Corona et al. (2019); Garetti and Taisch (2012). The CE concept has been around for decades, mostly attributed to Pearce and Turner (1989). Following a comprehensive study on its exact definition, Geissdoerfer et al. (2017) define CE as “a regenerative system in which resource input and waste, emission, and energy leakage are minimized by slowing, closing, and narrowing material and energy loops. This can be achieved through long-lasting design, maintenance, repair, reuse, remanufacturing, refurbishing, and recycling”. An important momentum of CE is that it is restorative or regenerative by intention and design Foundation (2013). In economies at scale, this typically necessitates implementing proper policies and legal frameworks that promote circularity, e.g., such as the Industry 5.0 framework of the European Commission.
@@
Van der Aalst et al. van der Aalst et al. (2023) find that the 10R framework of Reike et al. (2018) is particularly well-suited to support sustainable systems engineering. While indeed there seems to exist a good match due to the additional, refined R-imperatives of remanufacturing and recovering energy, the activities of R-frameworks are generally still too high-level to be actionable, and lack mechanisms for the formal assessment of sustainability under these strategies. A substantial amount of research has been done on reuse in the realm of digitized engineering domains, such as software Krueger (1992). Nowadays, it is hard to encounter top software conferences without dedicated tracks or focus topics on reuse, repurposing, or other kinds of value retention. As a consequence, the benefits Mikkonen and Taivalsaari (2019) and threats Gkortzis et al. (2021) of reuse are fairly understood in the software engineering community. The rich literature on the topic merits a deeper look from a systems engineering point of view to identify methods and techniques that can be reused to foster circular systems engineering.
@@ 3 Circular Systems Engineering
As put forward by van der Aalst et al. (2023), true sustainability requires circularity mechanisms. The idea of circularity has been around for decades and has been generally recognized as a desirable direction for humankind’s overall sustainability and innovation endeavors (Winans et al., 2017). While the state of the art considers circular economy Reike et al. (2018) as an enabler to sustainability, we argue that promoting circularity to a first principle in systems engineering fosters stronger alignment of technical and sustainability properties of engineered systems.
@@ 3.1 Circularity in systems engineering
Value retention in the definition refers to the act of keeping value as long as possible. Value retention is implemented through the mechanism of circularity. Circularity ensures physical and virtual value is reused, renewed, and regenerated, rather than being wasted. Waste is produced when value cannot be circulated anymore Morseletto (2020). While the governing frame of thinking contextualizes sustainability within the lifecycle of one specific system, circular systems engineering emancipates sustainability from these confines and unlocks a more holistic view of sustainability.
@@
Some extended R-frameworks (see Section 2.2) refine the reuse-recycle mechanisms into more tangible ones. For example, half of the R-imperatives of the 10R framework Reike et al. (2018) focus on retaining value by disassembly and reassembly—either with the same components, configuration, and purpose, or with completely new ones. Component traceability has been a topic of interest in Internet of Production (IoP) initiatives Jeschke et al. (2017); Pennekamp et al. (2019). In electrical and electronic equipment waste refurbishing services, for example, traceability through transportation is an important enabler of the offering Sharpe et al. (2018) as identifying transported products provides assurance of location, condition, and integrity. These early results should be further extended to traceability of components along multiple system lifetimes to contribute to more effective circularity mechanisms. Eventually, system components themselves are retired and disassembled, and in most cases, scrap raw material and energy are the only value that can be salvaged. Understanding material trade-offs, especially using better-choice materials—i.e., functional materials with reduced environmental impact by enhanced durability and lifetime Shen et al. (2023)—is paramount for reuse and repurposing. To assist such an understanding, machine learning and AI methods have been proposed, e.g., to accelerate material development by physics-constrained AI Gomes et al. (2019) and design space exploration by active transfer learning and data augmentation Kim et al. (2021). Such advanced mechanisms improve the effectiveness of value retention mechanisms in circular systems engineering and therefore, their development offers research avenues with elevated utility.
@@
Engineering knowledge encoded in digital and virtual value represents a particularly important class of values that should be retained across numerous system lifetimes. Yet, knowledge retention is not well-understood and is mostly ignored in current circularity strategies and frameworks Corona et al. (2019); Iacovidou et al. (2021). Efficient reuse of virtual value contributes to each aspect of sustainability as it allows for better design both in technical and economic terms, while also allowing for optimized operation and governance Proper et al. (2021); Poels et al. (2022). Reusing design knowledge and transposing previously learned lessons has been a topic of particular interest. Traditional techniques for knowledge retention through the use of ontologies provide rich expressiveness but do not scale to the level of truly complex systems and their engineering processes Osman et al. (2021). This is partly due to the lack of methods for identifying truly valuable knowledge to be retained. Assessing the value of knowledge is particularly challenging due to its non-monetary, highly abstract, and tacit nature. This is aptly exemplified by the discipline of infonomics Laney (2017), which takes a middle ground and provides actionable metrics for information, instead of knowledge. More automated techniques of retaining codified knowledge have been a subject of active research in machine learning. Transfer learning Weiss et al. (2016) is the technique of applying previously learned knowledge in congruent tasks. Such techniques have been successfully applied in an array of complex problems, such as atmospheric dust aerosol particle classification Ma et al. (2015) and poverty mapping Xie et al. (2016), and very much technical problems, such as image recognition Shao et al. (2015).
@@ 3.2 Impacting systems and methods through leverage points
Changing the goals of the systems engineering endeavor. Instead of optimizing for short-term, e.g., for the functional properties of one specific engineered system, operational excellence of organizations in circular systems engineering considers every engineered system and the goal of the organization becomes the overall sustainability of engineered systems and engineering methods.
@@
Discussions with technology transfer experts indicate that circular systems engineering might be an important development for industry domains, such as industrial automation Bangemann et al. (2014)–—the computer-aided technique of alleviating the human bottleneck in advanced industry processes. Driven by stakeholder demands for operational excellence Dahlgaard and Dahlgaard (1999) and better control over quality and time-to-market, industrial automation is a global market that is experiencing rapid growth and could be impacted at any of its L1—L3 leverage points by a circular systems engineering mindset.
@@ 4 Principles of Circular Systems Engineering
Circular systems engineering relies on two principles. First, fostering circularity mechanisms requires an understanding of how and when value is created along the systems engineering process. This is the principle of end-to-end sustainability (Section 4.1). Second, along with the sustainability of the engineered system, the sustainability of the employed engineering methods is equally important. This is the principle of bipartite sustainability (Section 4.2). Implementing circular systems engineering requires implementing both end-to-end and bipartite sustainability.
@@ 4.1 End-to-end sustainability
To achieve circularity in systems engineering, a detailed end-to-end understanding of the engineering endeavor is required. That is, reasoning about sustainability must scale from the scope of single engineering activities to the overall engineering process, and eventually, to networks of processes Oberdorf et al. (2023). In each scope, reasoning must entail $(i)$ assessment of sustainability and $(ii)$ finding optimal trade-offs among functional and sustainability properties.
@@
Assessing and optimizing for sustainability starts in the scope of single lifecycle phases or activities of the systems engineering process. This scope is typically implied by organizationally and logically isolated engineering silos. Engineering silos come in various sizes and complexity. For example, an engineering silo can be a team of hydraulics experts in the development process of a complex cyber-physical system, or a separate business unit or company integrated into the supply chain. Typical examples in this scope include reasoning for sustainability at design time Fatima and Lago (2023), reasoning for sustainability during manufacturing Henao-Hernández et al. (2019), and reasoning for sustainability at operation time Franciosi et al. (2018). Unfortunately, siloed methods do no longer scale with the pace of innovation and the expectations of quality and efficiency Dertien and Hastings (2021). Despite the high degree of digitalization, reasoning about sustainability within engineering silos has limitations due to multiple sources of truth and data inaccessibility, and challenges organizations by performance issues due to duplicated efforts.
@@
Advancements in digitalization and widespread digital transformation triggered organizations to place more focus on their business processes and foster end-to-end business process execution and management Eversheim (2013). In modern organizations, departments are connected through enterprise process networks Page and Opper (1999), linking departments, processes, and information systems and establishing inter-department and inter-process dependencies. Investigating sustainability in the scope of process networks allows for assessing sustainability and finding trade-offs by considering inter-process phenomena and properties. Figure 1 shows two typical cases in which parallel and sequential processes overlap and meet each other. One of the typical activities this scope enables is the rationalization of manufacturing. In Figure 1, the Manufacturing phases of Process 1 and Process 3 overlap. Synchronizing the manufacturing phases of engineering processes and optimizing production scheduling is one of the major energy-saving factors Gao et al. (2019) as about 80% of energy consumed by machine tools is reportedly attributed to idle state operation Mouzon et al. (2007). Optimizing the manufacturing schedule can substantially improve the energy footprint of organizations. This can be achieved by finding proper schedules across engineering processes. Another typical activity in this scope is design for reuse, i.e., allowing future generations to retain value from a retired system more easily. As the post-life phase of one engineering process is channeled into the concept design phase of a subsequent process, substantial value can be retained in the form of raw material, energy, and engineering knowledge. The scope of process networks is where circularity emerges as value is being circulated from one engineered system to the other. This allows for a higher level of sustainability, specifically in environmental and economic aspects. Fostering circular systems engineering, thus, requires an end-to-end understanding of process networks.
@@ 4.2 Bipartite sustainability
True sustainability can only be achieved if sustainable systems are engineered through sustainable methods. We refer to this principle as bipartite sustainability. Bipartite sustainability emphasizes that sustainable systems and sustainable engineering methods are equally important and form an inseparable unit when assessing sustainability, as incurring sustainability debts along the engineering process might easily render the overall, end-to-end systems engineering endeavor unsustainable.
@@
Another supporting method of sustainability ambitions that itself requires sustainability considerations is digital twinning (Kritzinger et al., 2018). Capgemini reports that 60% of organizations believe twin technology is critical to improving sustainability efforts Capgemini (2022). Some of the sectors with visible benefits include manufacturing, where multi-million dollar savings are being regularly reported thanks to digital twinning The Wall Street Journal Online (2019); smart cities, where Accenture estimates energy consumption rationalization by 30%-80% through digital technologies Accenture (2022); and traditionally lower-digitized sectors, such as biophysical systems and agriculture, where digitalization improves crop-to-energy ratio and substantially stabilizes the supply chain David et al. (2023a). While research mostly focuses on sustainability by digital twins, it is important to acknowledge that digital twins themselves need to be sustainable to support larger sustainability goals. Bellis and Denil (2022) report four important sustainability challenges of digital twinning: energy consumption, modeling effort and complexity, the ability to evolve with the physical twin, and the deployment of the twin architecture within organizations. The development of digital twins is hindered by the complexity of systems subject to digital twinning. To improve the economic sustainability of digital twins, component-based techniques have been recommended, e.g., through component-based techniques. Variability has been recommended to improve the composability of digital twins Berger et al. (2013); Kang and Lee (2013), but foundational composition-based techniques are lacking. Closely related work has been done in the synthesis of simulators Hansen et al. (2021), which are key enablers of digital twins. Such preliminary results could be extended to the combined digital-physical realm to support the rapid synthesis of digital twins. From a technical sustainability point of view, proper evolution mechanisms of digital twins, potentially governed by comprehensive frameworks and taxonomies David and Bork (2023) are an excellent way toward more sustainable digital methods. On a related note, self-adaptation has been identified by Fur et al. (2023) as a key mechanism to technical sustainability of digital twins that require, e.g., flexible architectures, automated re-calibration of simulation models, and ontologies to augment structural information.

### 20 Jan 2024 v3

@@ 1 Introduction
The steadily accelerating innovation pathways of humankind have rendered prevailing systems engineering paradigms unsustainable. By Brundtland’s classic definition of sustainability (Brundtland, 1987), systems engineering falls short of “meeting the needs of the present without compromising the ability of future generations to meet their own needs”. By the terms of the four essential sustainability dimensions of Penzenstadler and Femmer (2013), prevalent systems engineering practices fail to fulfill important technical (long-term usage), economic (financial viability), environmental (reduced impact), and social (elevated utility) sustainability principles.
@@
Clearly, systems engineering must change course. While most experts agree that effective change will require major political and industry intervention Freitag et al. (2021), and some argue that fulfilling these goals is not feasible due to the limitations of our current frame of thinking Becker (2023), it is also expected that, in the decade ahead of us, users and organizations will reward and demand efforts toward sustainability. The importance of reuse and repurposing will increase and become a key driving force in systems engineering. Sustainability as a characteristic will become a leading principle in the design, operation, maintenance, and post-life of systems. These trends have been collectively identified in the Systems Engineering Vision 2035 report International Council on Systems Engineering (2021) (INCOSE) of the International Council on Systems Engineering (INCOSE) as the top “global megatrend” that will instigate the development of radically new systems engineering frameworks, methods, and tools. The European Commission has also identified sustainability as a critical enabler of a more resilient European industry within the framework of Industry 5.0 European Commission (2020). Expert voices call for immediate action in devising frameworks, methods, and tools for sustainable systems engineering practices van der Aalst et al. (2023) and fostering a circular economy Kristoffersen et al. (2020).
@@
The rest of this article is structured as follows. In Section 2, we provide an overview of the background related to our work. In Section 3, we offer a definition of circular systems engineering. In Section 4, we discuss the foundational principles of circular systems engineering. In Section 5, we discuss sustainability maturity levels and strategies to reach the level of circularity in systems engineering. In Section 6, we outline the key challenges and research opportunities for the modeling community in implementing circular systems engineering. In Section 7, we draw the conclusions.
@@ 2.2 Circular economy
Circular economy (CE) is a model of production and consumption in an economic system that promotes retaining value over numerous cycles of the system as long as possible. CE is one among several solutions for fostering sustainable systems Corona et al. (2019); Garetti and Taisch (2012). The CE concept has been around for decades, mostly attributed to Pearce and Turner (1989). Following a comprehensive study on its exact definition, Geissdoerfer et al. (2017) define CE as “a regenerative system in which resource input and waste, emission, and energy leakage are minimized by slowing, closing, and narrowing material and energy loops. This can be achieved through long-lasting design, maintenance, repair, reuse, remanufacturing, refurbishing, and recycling”. An important momentum of CE is that it is restorative or regenerative by intention and design Foundation (2013). In economies at scale, this typically necessitates implementing proper policies and legal frameworks that promote circularity, such as the Industry 5.0 framework of the European Commission.
@@
Van der Aalst et al. van der Aalst et al. (2023) find that the 10R framework of Reike et al. (2018) is particularly well-suited to support sustainable systems engineering. Although with the additional refined R-imperatives 10R indeed provides a proper coverage of systems engineering concerns, the activities of R-frameworks are generally too high-level to be actionable, and lack mechanisms for the formal assessment of sustainability under these strategies. A substantial amount of research has been done on reuse in the realm of digitized engineering domains, such as software Krueger (1992). Nowadays, it is hard to encounter top software conferences without dedicated tracks or focus topics on reuse, repurposing, or other kinds of value retention. As a consequence, the benefits Mikkonen and Taivalsaari (2019) and threats Gkortzis et al. (2021) of reuse are fairly understood in the software engineering community. The rich literature on the topic merits a deeper look from a systems engineering point of view to identify methods and techniques that can be reused to foster circular systems engineering.
@@ 3 Circular Systems Engineering
As put forward by van der Aalst et al. (2023), true sustainability requires circularity mechanisms. The idea of circularity has been around for decades and is generally recognized as a desirable direction for humankind’s sustainability and innovation endeavors (Winans et al., 2017). While the state of the art considers circular economy Reike et al. (2018) as an enabler of sustainability, we argue that promoting circularity to a first principle in systems engineering fosters stronger alignment of technical and sustainability properties of engineered systems.
@@ 3.1 Circularity in systems engineering
In the definition, value retention refers to the act of keeping value as long as possible. Value retention is implemented through the mechanism of circularity. Circularity ensures physical and virtual value is reused, renewed, and regenerated, rather than being wasted. Waste is produced when value cannot be circulated anymore Morseletto (2020). While the governing frame of thinking contextualizes sustainability within the lifecycle of one specific system, circular systems engineering emancipates sustainability from these confines and unlocks a more holistic view of sustainability.
@@
Some extended R-frameworks (see Section 2.2) refine the reuse-recycle mechanisms into more tangible ones. For example, half of the R-imperatives of the 10R framework Reike et al. (2018) focus on retaining value by disassembly and reassembly—either with the same components, configuration, and purpose, or with completely new ones. Component traceability has been a topic of interest in Internet of Production (IoP) initiatives Jeschke et al. (2017); Pennekamp et al. (2019). In electrical and electronic equipment waste refurbishing services, for example, traceability through transportation is an important enabler of the offering Sharpe et al. (2018) as identifying transported products provides assurance of location, condition, and integrity. These early results should be further extended to the traceability of components along multiple system lifetimes to contribute to more effective circularity mechanisms. Eventually, system components are retired and disassembled, and in most cases, scrap raw material and energy are the only value that can be salvaged. Understanding material trade-offs, especially using better-choice materials—i.e., functional materials with reduced environmental impact by enhanced durability and lifetime Shen et al. (2023)—is paramount for reuse and repurposing. To assist such an understanding, machine learning and AI methods have been proposed, e.g., to accelerate material development by physics-constrained AI Gomes et al. (2019) and design space exploration by active transfer learning and data augmentation Kim et al. (2021). Such advanced mechanisms improve value retention in circular systems engineering, and therefore, their development offers research avenues with elevated utility.
@@
Engineering knowledge encoded in digital and virtual value represents a particularly important class of values that should be retained across numerous system lifetimes. Yet, knowledge retention is not well-understood and is mostly ignored in current circularity strategies and frameworks Corona et al. (2019); Iacovidou et al. (2021). Efficient reuse of virtual value contributes to each aspect of sustainability as it allows for better design both in technical and economic terms, while also allowing for optimized operation and governance Proper et al. (2021); Poels et al. (2022). Reusing design knowledge and transposing previously learned lessons has been a topic of particular interest. Traditional techniques for knowledge retention through the use of ontologies provide rich expressiveness but do not scale to the level of truly complex systems and their engineering processes Osman et al. (2021). This is partly due to the lack of methods for identifying truly valuable knowledge to be retained. Assessing the value of knowledge is particularly challenging due to its non-monetary, highly abstract, and tacit nature. This is aptly exemplified by the discipline of infonomics Laney (2017), which takes a middle ground and provides actionable metrics for information instead of knowledge. More automated techniques of retaining codified knowledge have been a subject of active research in machine learning. Transfer learning Weiss et al. (2016) is the technique of applying previously learned knowledge in congruent tasks. Such techniques have been successfully applied in an array of complex problems, such as atmospheric dust aerosol particle classification Ma et al. (2015) and poverty mapping Xie et al. (2016), and very much technical problems, such as image recognition Shao et al. (2015).
@@ 3.2 Impacting systems and methods through leverage points
Changing the goals of the systems engineering endeavor. Instead of optimizing for short-term, e.g., for the functional properties of one specific engineered system, operational excellence of organizations in circular systems engineering considers every engineered system, and the goal of the organization becomes the overall sustainability of engineered systems and engineering methods.
@@
Discussions with technology transfer experts indicate that circular systems engineering might be an important development for industry domains, such as industrial automation Bangemann et al. (2014)–—the computer-aided technique of alleviating the human bottleneck in advanced industry processes. Driven by stakeholder demands for operational excellence Dahlgaard and Dahlgaard (1999) and better control over quality and time-to-market, industrial automation is a global market experiencing rapid growth, and could be impacted at any of its L1—L3 leverage points by a circular systems engineering mindset.
@@ 4 Principles of Circular Systems Engineering
Circular systems engineering relies on two principles. First, fostering circularity mechanisms requires understanding how and when value is created along the systems engineering process. This is the principle of end-to-end sustainability (Section 4.1). Second, along with the sustainability of the engineered system, the sustainability of the employed engineering methods is equally important. This is the principle of bipartite sustainability (Section 4.2). Implementing circular systems engineering requires implementing both end-to-end and bipartite sustainability.
@@ 4.1 End-to-end sustainability
To achieve circularity in systems engineering, a detailed end-to-end understanding of the engineering endeavor is required. That is, reasoning about sustainability must scale from the scope of single engineering activities to the overall engineering process and, eventually, to networks of processes Oberdorf et al. (2023). In each scope, reasoning must entail $(i)$ assessment of sustainability and $(ii)$ finding optimal trade-offs among functional and sustainability properties.
@@
Assessing and optimizing for sustainability starts in the scope of single lifecycle phases or activities of the systems engineering process. This scope is typically implied by organizationally and logically isolated engineering silos. Engineering silos come in various sizes and complexity. For example, an engineering silo can be a team of hydraulics experts in the development process of a complex cyber-physical system, or a separate business unit or company integrated into the supply chain. Typical examples in this scope include reasoning for sustainability at design time Fatima and Lago (2023), reasoning for sustainability during manufacturing Henao-Hernández et al. (2019), and reasoning for sustainability at operation time Franciosi et al. (2018). Unfortunately, siloed methods no longer scale with the pace of innovation and the expectations of quality and efficiency Dertien and Hastings (2021). Despite the high degree of digitalization, reasoning about sustainability within engineering silos has limitations due to multiple sources of truth and data inaccessibility, and challenges organizations by performance issues due to duplicated efforts.
@@
Advancements in digitalization and widespread digital transformation triggered organizations to focus more on their business processes and foster end-to-end business process execution and management Eversheim (2013). Modern organizations connect departments through enterprise process networks Page and Opper (1999), linking departments, processes, and information systems and establishing inter-department and inter-process dependencies. Investigating sustainability in the scope of process networks allows for assessing sustainability and finding trade-offs by considering inter-process phenomena and properties. Figure 1 shows two typical cases in which parallel and sequential processes overlap and meet each other. One of the typical activities this scope enables is the rationalization of manufacturing. In Figure 1, the Manufacturing phases of Process 1 and Process 3 overlap. Synchronizing the manufacturing phases of engineering processes and optimizing production scheduling are major energy-saving factors Gao et al. (2019) as about 80% of the energy consumed by machine tools is reportedly attributed to idle state operation Mouzon et al. (2007). Optimizing the manufacturing schedule can substantially improve the energy footprint of organizations. This can be achieved by finding proper schedules across engineering processes. Another typical activity in this scope is design for reuse, i.e., allowing future generations to retain value from a retired system more efficiently. As the post-life phase of one engineering process is channeled into the concept design phase of a subsequent process, substantial value can be retained in the form of raw material, energy, and engineering knowledge. The scope of process networks is where circularity emerges as value is circulated from one engineered system to the other. This allows for a higher level of sustainability, specifically in environmental and economic aspects. Fostering circular systems engineering, thus, requires an end-to-end understanding of process networks.
@@ 4.2 Bipartite sustainability
True sustainability can only be achieved if sustainable systems are engineered through sustainable methods. We refer to this principle as bipartite sustainability. Bipartite sustainability emphasizes that sustainable systems and sustainable engineering methods are equally important and form an inseparable unit when assessing sustainability, as incurring sustainability debts along the engineering process might render the overall, end-to-end systems engineering endeavor unsustainable.
@@
Another supporting method of sustainability ambitions that itself requires sustainability considerations is digital twinning (Kritzinger et al., 2018). Capgemini reports that 60% of organizations believe twin technology is critical to improving sustainability efforts Capgemini (2022). Some of the sectors with visible benefits include manufacturing, where multi-million dollar savings are being regularly reported thanks to digital twinning The Wall Street Journal Online (2019); smart cities, where Accenture estimates energy consumption rationalization by 30%-80% through digital technologies Accenture (2022); and traditionally lower-digitized sectors, such as biophysical systems and agriculture, where digitalization improves crop-to-energy ratio and substantially stabilizes the supply chain David et al. (2023a). While research primarly focuses on sustainability by digital twins, it is important to acknowledge that digital twins themselves need to be sustainable to support larger sustainability goals. Bellis and Denil (2022) report four sustainability challenges of digital twinning: energy consumption, modeling effort and complexity, the ability to evolve with the physical twin, and deploying the twin architecture within organizations. The development of digital twins is hindered by the complexity of systems subject to digital twinning. To improve the economic sustainability of digital twins, variability-based techniques have been recommended Berger et al. (2013); Kang and Lee (2013), but sound composition mechanisms are still lacking. Closely related work has been done in the synthesis of simulators Hansen et al. (2021), which are key enablers of digital twins. Such preliminary results could be extended to the combined digital-physical realm to support the rapid synthesis of digital twins. From a technical sustainability point of view, proper evolution mechanisms of digital twins, potentially governed by comprehensive frameworks and taxonomies David and Bork (2023), are an excellent way to move toward more sustainable digital methods. On a related note, self-adaptation has been identified by Fur et al. (2023) as a key mechanism for the technical sustainability of digital twins that require, e.g., flexible architectures, automated re-calibration of simulation models, and ontologies to augment structural information.